OpenStudy (anonymous):

i am stuck on this question can someone plz help me? rationalize the denominator of 5/(6+3sqrt2)-(2+2sqrt2) A.20-5sqrt2/14 B.20+5sqrt2/18 C.20-sqrt2/14 D.5sqrt2/2+4sqrt2

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

@damien

OpenStudy (anonymous):

(6+3sqrt2)-(2+2sqrt2) the denominator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Simplify the denominator first by collecting the like terms.. (sorry I was late, I forgot about this question)

OpenStudy (jdoe0001):

$$\bf \frac{5}{(6+3\sqrt{2})-(2+2\sqrt{2})}\\ \color{blue}{\text{multiplying both sides by the denominator's conjugate}}\\ \frac{5}{(6+3\sqrt{2})-(2+2\sqrt{2})} \times \frac{(6+3\sqrt{2})+(2+2\sqrt{2})}{(6+3\sqrt{2})+(2+2\sqrt{2})}\\ \color{blue}{\text{keep in mind that } \implies (a-b)(a+b) = (a^2-b^2)}\\ \frac{5\times [(6+3\sqrt{2})+(2+2\sqrt{2})]}{(6+3\sqrt{2})^2-(2+2\sqrt{2})^2} $$

OpenStudy (anonymous):

\[\frac{ 5 }{ \left( 6+3\sqrt{2} \right)-\left( 2+\sqrt{2} \right) }\] \[\frac{ 5 }{6+3\sqrt{2}-2-2\sqrt{2} }=\frac{ 5 }{4+1\sqrt{2} }\] \[\frac{ 5 }{ 4+\sqrt{2}} \times \frac{ 4-\sqrt{2} }{4-\sqrt{2}}=\frac{ 5\left( 4-\sqrt{2} \right) }{ 4^{2}-\left( \sqrt{2 }\right)^2 }\] \[=\frac{ 5\left( 4-\sqrt{2} \right) }{16-2 }=\frac{ 20-5\sqrt{2} }{ 14 }\]

OpenStudy (anonymous):

thanks so much for your help i understand better now